OpenStudy (anonymous):

determine if h(x)=5/x+2 is an inverse of f(x)=5/x-2

OpenStudy (anonymous):

The definition of an inverse: \[ f((h(x))=h(f(x))=x \]So, does: \[ \frac{5}{f(x)+2}=\frac{5}{\frac{5}{x-2}+2}=x? \]